OpenStudy (anonymous):

I am working on reducing ratios to lowest terms. The problem I came across was reducing 14.2 : 71. The answer the book gave was 1 : 5. Can someone please tell me how they came to this answer?

OpenStudy (anonymous):

Replace the : symbol with / and then simplify. 14.2:71 -> 14.2/71= 0.2 = 2/10 =1/5= 1:5

OpenStudy (anonymous):

That was easy. Thank you very much.

OpenStudy (anonymous):

Have a good day.